OpenStudy (anonymous):

the annual interest rate on a $6000 investment exceeds the interest earned on a $4000 investment by $204. The $6000 is invested at a 0.4% higher rate of interest than the $4000. What is the interest rate of each investment?

OpenStudy (anonymous):

Assuming interest earned in a year, r(1) the rate on the $6000 investment, r(2) the rate on the $4000 investment, we have 6000(1 + r(1)) = 4000(1 + r(2)) + 204, r(1) = r(2) + .004. Now that you have two equations in two unknowns, you can solve for them. The system, as is, can be solved by the substitution method.

OpenStudy (anonymous):

I still don't understand!